Find the domain of the function:
f(x) = 5-3x/2x^2-14x+20

Answer by lwsshak3(11628)

Find the domain of the function:
f(x) = 5-3x/2x^2-14x+20
set denominator=0, then solve for x
At these x-values f(x), becomes undefined and therefore they are not included in the domain
..
2x^2-14x+20=0
2(x^2-7x+10)=0
2(x-5)(x-2)=0
x=2 and 5
Domain: x≠2 and x≠5
or in interval notation
(-∞,2) U (2,5) U (5,∞)